OpenStudy (anonymous):

Write and simplify the equation you would use to solve the problem below. A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A. w + 5 = 116 B. 4w + 10 = 116 C. 2w + 5 = 116 D. 2w + 10 = 116

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

@blurbendy

OpenStudy (johnweldon1993):

Well we know that the perimeter is the sum of all the sides so \[P = 2w + 2l\] And we are given to understand that the length is 5 feet more than the width so \[l = w + 5\] And we would substitute that in for 'l' in our first equation...can you do that?

OpenStudy (johnweldon1993):

actually we can plug in the value for perimeter as well since we know it our 2 equations are \[116 = 2w + 2l\] \[l = w + 5\]

OpenStudy (johnweldon1993):

So NOW substitute that stuff we found out that l equals.....and plug that into our 1st equation \[116 = 2w + 2(w + 5)\] Can you solve it from here?